# Emission and Absorption Lines in Photospheric Radius Expansion Bursts of 4U 1820–30

### v1

@@ 4 Conclusions
In this study, we carried out a comprehensive spectral analysis of 4U 1820–30 using NICER observations, focusing on the search for emission and absorption lines in 15 PRE bursts. We identified a 1 keV emission line from 14 bursts, a 1.6 keV absorption line from one burst, and a 3 keV absorption line from 12 bursts. The main results are generally consistent with those derived from the first five bursts presented by Strohmayer et al. (2019); however, our findings indicate a lower confidence level for the 1.6 keV absorption line. These lines were detected in coadded spectra, revealing a significant and systematic shift in their energies between bursts with varying PRE radii. These shifts are consistent with predictions from burst-driven wind models, suggesting that the spectral lines originate from heavy elements in the PRE winds. Based on the predicted line energies in winds, we attribute the 1 keV emission line to Fe L-shell transitions. Furthermore, our analysis indicates that the 1 keV line is a superposition of several narrow lines rather than a single broad feature. Simulated mock observations with a 1-second exposure time demonstrate that HUBS data can effectively resolve the 1 keV emission lines during bursts, offering new opportunities for precise spectral studies.

### v2

@@ 4 Conclusions
In this study, we carried out a comprehensive spectral analysis of 4U 1820–30 using NICER observations, focusing on the search for emission and absorption lines in 15 PRE bursts. We identified a 1 keV emission line from 14 bursts, a 1.6 keV absorption line from one burst, and a 3 keV absorption line from 12 bursts. The main results are generally consistent with those derived from the first five bursts presented by Strohmayer et al. (2019). However, our findings indicate a lower confidence level for the 1.6 keV absorption line. These lines were detected in coadded spectra, revealing a significant and systematic shift in their energies between bursts with varying PRE radii. These shifts are consistent with predictions from burst-driven wind models, suggesting that the spectral lines originate from heavy elements in the PRE winds. Based on the predicted line energies in winds, we attribute the 1 keV emission line to Fe L-shell transitions. Furthermore, our analysis indicates that the 1 keV line is a superposition of several narrow lines rather than a single broad feature. Simulated mock observations with a one-second exposure time demonstrate that HUBS data can effectively resolve the 1 keV emission lines during bursts, offering new opportunities for precise spectral studies.